OpenStudy (alleeyyssaa):

help with math please!!!! Statistical models predict that the price (in dollars) of a 500 GB hard drive will change according to the function p(t) = 600 − 2t2, where t is the month. Which expression gives the number of months t (passed since January 1) in terms of the price p?

OpenStudy (anonymous):

ok i ll give u hint then we ll work from there do u know how to make something the subject e.g. y=x+2 and i wanna make x the subject i ll write x= y-2 . so in this question u must make t the subject.

OpenStudy (alleeyyssaa):

i have no idea because i dont have an actual teacher to teach me this stuff

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

but did u understood what i said on top

OpenStudy (alleeyyssaa):

a little but not very well

OpenStudy (anonymous):

ok so this i s ur expression \[p(t)=600-2t^2\]

OpenStudy (alleeyyssaa):

okay

OpenStudy (anonymous):

now u have to manipulate this expression such that it looks something like "t=....."

OpenStudy (alleeyyssaa):

would it be like t=600-2p^2

OpenStudy (alleeyyssaa):

or no?

OpenStudy (anonymous):

no

OpenStudy (alleeyyssaa):

shoot, okay so what i do then?

OpenStudy (anonymous):

ok first take 2t^2 to other side and bring p(t) to other

OpenStudy (anonymous):

try it , sw if its wrong